OpenStudy (candycane):

Hello I am in Algebra 1 takening the 2nd half this school year with Connections academy, I am working on some extra credit points. Right now Im working on the quadratic functions and equations practice #4. Find the axis of symmetry for y=5x^2-7x+2. Could someone please explain how to come to the correct answer?? Thanks!

OpenStudy (anonymous):

half way between the x intercepts is the axis of symmery

OpenStudy (candycane):

Okay I imput the equation into a graphing calculator and the answers came out to be 0.7,-0.45

OpenStudy (anonymous):

I get x = 7/10

OpenStudy (anonymous):

not sure where you're getting the -.45

OpenStudy (candycane):

Thats what the graphing calc gave me. The test says x= 7/10 i think the -.45 is the y part of it Im not sure

OpenStudy (ranga):

This is a parabola. You can complete the square and find the x-axis of the vertex and a vertical line through the vertex will be the axis of symmetry. It is x = -b/2a = 7/10

OpenStudy (candycane):

okay that sorta makes sense

OpenStudy (ranga):

Yes the vertex is at (0.7, -0.45)

OpenStudy (candycane):

but it only let me choose one, so the axis of symmetry would be x and x would be 7/10

OpenStudy (ranga):

The equation for the axis of symmetry is x = 7/10

OpenStudy (candycane):

okay that makes sense, i think i've got it

OpenStudy (candycane):

Thanks everyone :3

OpenStudy (anonymous):

Your welcome.